# Machine Learning Framework for Characterizing Processing–Structure Relationship in Block Copolymer Thin Films

### v1

@@ 2.2.2 AFM
The BCP thin films in this study are expected to arrange in hexagonal cylinder morphologies, given the PS-b-PEO block ratio. However, the orientation and ordering of the cylinders is highly dependent on the processing conditions of the BCP films. 59 Therefore, the 202 AFM images collected represent various orientations of hexagonally packed cylinders, including vertical and horizontal alignments, which manifest as ‘dot’- and ‘line’-type features in the AFM images, respectively. While no single approach is sufficient to analyze the local ordering, or grain size, of the microdomains for different alignments, multiple approaches can be used in tandem to create a robust AFM analysis approach (Fig. 3).
@@
Figure 3: Overview of the high-throughput AFM image analysis process for extracting grain size. a) AFM images are classified based on morphology by a convolutional neural network (CNN). b) Images of mixed-type morphology are segmented into dot- and line-type features. c) Color-wheel analysis of line-type images, color-coding microdomains based on orientation, then segmenting the color-coded microdomains before determining grains. d) Voronoi analysis of dot-type images, features are first isolated before applying the Voronoi analysis and identifying grains.
@@
AFM images are first classified into three categories (Fig. 3 a) depending on the alignment of the hexagonal cylinders with the use of a convolutional neural network (CNN), discussed in more detail in the following section. After classification, the binarized dot- and line-type images were ready for analysis, however, the mixed-type images required further pre-processing. Segmentation of the mixed-type images was performed using the labelme software to create two distinct masks for the dot- and line-type features (Fig. 3 b). 60 The segmented masks were treated the same as the the dot- and line-type images for further analysis.
For line-type images, the MLExchange color wheel was applied to the Two-dimensional fast Fourier transforms (2D FFT) of the AFM images, which assigns a unique color to the microdomains of different orientations (Fig. 3 c).61 After isolating one phase from the color wheel image, i.e. filtering out one of the blocks, the microdomains were segmented by creating masks related to each orientation. Once isolated, the scikit-learn cluster package was used to group adjacent pixels together, which were then filtered according to a size threshold to reduce noise in the images.62 Individual grains were identified by another clustering analysis before applying another threshold to filter out grains smaller than 150% of the average, again to reduce noise and artificial grains. The final grain size was then calculated by averaging the grain sizes for each sample, then converting from pixels2 to $\mu$m2. In total, this grain analysis process takes approximately 8 seconds per image. Furthermore, the in-house grain analysis code was evaluated, and results were generally in agreement with grain analysis from a separate study63 which is discussed in more detail in the SI.
Separately, the dot-type images were analyzed using Voronoi analysis (Fig. 3 d). Traditionally, Voronoi type analyses are used to identify grain boundaries and defects in well-ordered BCP systems.64, 65 The following approach, however, leverages the Voronoi analysis to characterize two-dimensional grain sizes for weakly-ordered systems. Within the analysis, the dot-type features were first isolated using the labelme software, creating discrete points for each dot-type feature. Voronoi cells were constructed by treating each feature in the image as a center point. The image was divided such that each cell contains the area closer to its feature than to any other, resulting in a set of perfectly adjacent, non-overlapping regions—one surrounding each dot-like feature. For an ideal hexagonal cylinder morphology, each Voronoi cell should have 6 equal edges, reflecting 6 surrounding hexagonally packed cylinders. Therefore, the numbers of edges for each cell were determined, additionally having to satisfy a criteria being within a 55% threshold of the mode edge length in the image, in order to filter edges at the end of the image or near a defect. Additionally, for a cell to be considered for the grain analysis, the 6 edges must be within 2.5% of each other, ensuring well-ordered hexagonal packing. Finally, adjacent cells were grouped together, and the average grain size determined across all grains for a given sample. The Voronoi analysis occurs rapidly, taking only 3 seconds per image.
@@
where $u$ is the peak position of the radial distribution, and $FOV$ is the field of view of the AFM images (i.e. image size, such as 1 $\mu$m2 for images in Fig. 3).66 This process occurs rapidly, taking less than 10 seconds for all 202 samples, while maintaining a high degree of fitting accuracy with an $R^{2}$ of 0.96.
@@ 2.2.3 Convolutional Neural Network AFM Image Classification
A CNN was first employed to classify the AFM images based on their morphological alignments, specifically, line-, dot-, and mixed-type features, corresponding to vertically aligned cylinders, horizontally aligned hexagonal cylinders, and a combination of both. The CNN was trained using the binarized images as inputs and manually labeled classes as targets. Additionally, due to a significant imbalance of images with line-type morphologies relative to the images with dot- and mixed-type morphologies, the images within the latter classes were augmented, creating three copies of each image rotated in increments of 90°, creating a more balanced dataset. Furthermore, the dot-type images were reflected horizontally and vertically to further increase the amount of data for the class. After augmentation, the dataset was much more balanced, with 171, 84, and 72 images for line-, dot-, and mixed-type morphologies, respectively.
@@
The model follows a simplified VGG-like structure,67 consisting of three convolutional blocks with two convolutional layers each, followed by batch normalization, ReLU activations, and max pooling (Fig. 4 a). A more detailed description of the CNN can be found in the SI. Among the 327 AFM images used in the classification task, 228 were allocated for training, 65 for validation, and 34 for testing. On the test set, the CNN model achieved a high classification accuracy of 97.06%. To interpret the model’s decision-making process, the Shapley Additive exPlanations (SHAP) Python package was employed (Fig. 4 b-d). SHAP analysis provides a unified framework for explaining ML model predictions by assigning each feature an importance value—its SHAP value—based on its contribution to the model’s output.68
@@
Figure 4: a) Simplified architecture of the CNN used for morphological classification. AFM images are passed through a series of convolutional blocks to extract hierarchical feature maps, which are then fed into fully connected layers to produce a probabilistic prediction over the three morphological classes (line, dot, and mixed). b-d) SHAP-based interpretation of the CNN’s predictions at convolutional block 1-3. Positive SHAP values (red) highlight features that positively contribute to the predicted class, while negative SHAP values (blue) indicate features that negatively influence the prediction. e) Confusion matrix summarizing the CNN’s performance on the test set.
@@
Within each convolutional block, features were extracted from the input to generate feature maps. After each block, spatial dimensions were reduced, causing fine-scale features to be captured in the earlier blocks, while progressively larger and more abstract features were analyzed in subsequent blocks. The SHAP analysis of block 1 reveals that features evaluated at smaller length scales exhibit strong positive correlations with dot- and mixed-type morphologies, such as dots and disconnections in the line features. This aligns with the physical interpretation of the images, where these morphologies characteristically lack prominent larger structural features. Similarly, the SHAP analysis of block 2 reveals positive correlation between clusters of dots with the dot-type morphology, and groups and lines with heavy curvature with the mixed-type morphologies. Interestingly, block 2 also begins to pick up parallel groups of line features, positively correlating them with the line-type morphology. Finally, features extracted from block 3 primarily correlate positively with the line-type morphology, focusing on long stretches of line features. Overall, the CNN performed exceptionally well distinguishing between dot- and line-type morphologies, however, struggled with the mixed-type morphologies, over-predicting the dot-type morphology for some mixed-type images (Fig. 4 e).
@@ 3.1 Dataset overview
As mentioned previously, the processing parameters explored in this study include solvent ratios of toluene:THF from 0.5 to 1.0, additive ratios from 1.00 to 7.00, and additive types of CN and MN. Domain spacing and FWHM from GISAXS measurements show narrow distributions, with domain spacing measurements around 36 nm and FWHM around 0.0025 A-1 (Fig. 5 a, b). These distributions indicate the majority of the BCP films have relatively good ordering, indicated by the low FWHM value which is reflective of a narrow distribution of domain spacings in the sample. However, the AFM-measured morphological properties, domain spacing and grain size, demonstrate broader distributions (Fig. 5 c, d). While the range of values for the GISAXS and AFM-based measurements are in agreement with each other, the localized nature of the AFM measurements leads to higher variability, in contrast to GISAXS measurements which reflect the bulk properties of the sample.
@@
Pearson correlation coefficients were calculated to assess potential linear relationships between the processing parameters and the resulting morphological properties (Fig. 5 e). Among all processing parameters, the additive ratio exhibited the strongest Pearson correlation with the measured morphological properties. A previous study using in situ ellipsometry showed that it takes a few minutes for the high boiling point additive to evaporate from the BCP films.69 During this period, the additive-infused BCP film exhibits enhanced segmental mobility, allowing polymer chains to rearrange into more ordered and thermodynamically favorable morphologies. Interestingly, the type of additive exhibits a relatively strong Pearson correlation with both the GISAXS-measured FWHM and AFM-measured domain spacing. For AFM, this may be attributed to surface-specific effects introduced by the additive, such as differential evaporation dynamics, not captured by bulk GISAXS measurements. In the case of GISAXS-measured FWHM, the correlation may reflect broader distributions of domain spacings within the sample, suggesting increased structural heterogeneity without significantly altering the average domain spacing. Meanwhile, solvent ratio has a relatively low Pearson correlation across all measured outputs. Although solvent plays a critical role in the initial solubilization of the polymers, it has little effect on the resulting morphology of the BCP film, due to the low boiling point solvent volatilizing during the spin coating process. Furthermore, the Pearson correlations were evaluated for the morphological outputs with respect to each other (Fig. 5 f). The domain spacing from both GISAXS and AFM has a fairly large Pearson correlation, indicating good agreement between the measurements from the two approaches. Domain spacing from GISAXS and AFM are inversely correlated with GISAXS-measured FWHM, indicating that as the samples become more ordered, i.e., decreasing FWHM, domain spacing increases. Additionally, the grain size has positive Pearson correlations with domain spacing and an inverse correlation with FWHM, again indicating that as the samples become more ordered, the grain size increases. Scatter plots of the morphological properties are shown in the SI.
3.2 Processing–structure classification
3.2.1 Processing condition-trained model
@@
Classification models were trained using solvent ratio, additive ratio, and additive type as input features, with manually labeled ‘weakly-ordered’ and ‘well-ordered’ morphologies - based on AFM images - as the target output. Support Vector Classifier (SVC) and Random Forest Classifier (RFC) were selected for this task due to their robust performance and high accuracy in classification problems. The data set was split using 80% as the training set, and 20% for the test set to evaluate the model performance while using a 5-fold cross validation to ensure generalizability. A grid search approach was employed to evaluate various combinations of hyperparameters, which can be found in the SI for each model. The PyTorch accuracy metric was used to evaluate the models, in which the model with the best cross-validation accuracy was selected.
Both classifiers performed exceedingly well, with RFC performing slightly better than SVC, with cross-validation accuracies of 0.98 and 0.95, respectively. Notably, the RFC model over-predicted the ‘well-ordered’ class, likely due to the class imbalance, with the majority of data falling within the ‘well-ordered’ category (Fig. 6 a). Given the prevalence of SVC in classification tasks, the results were initially surprising. However, due to the relatively small amount of data, RFC was likely able to effectively map the classification task through a series of decision trees, resulting in only three incorrect classifications at the boundary (Fig. 6 b). Therefore, with a relatively small dataset RFC is highly capable of classification tasks, however, SVC could outperform RFC as the dataset increases in both size and complexity.
Figure 6: a) Confusion matrix showing the testing results of the RFC model. b) Classification plots illustrating the transition from ‘weakly-ordered’ to ‘well-ordered’ samples as a function of solvent and additive ratios. The plots are separated into two panels: top for the use of CN as the additive and bottom for the use of MN as the additive. Incorrect predictions by the RFC are marked with a red ‘x’.
3.2.2 AFM image-trained model
Further utilizing the AFM images, an additional CNN was trained using the binarized images as inputs and the manually labeled classes of ‘weakly-ordered’ and ‘well-ordered’ as outputs. Additionally, due to an imbalance of well-ordered to weakly-ordered images, 161 and 40 respectively, the images within the weakly-ordered class were augmented, creating three copies of each image rotated in increments of 90 °, creating a more balanced dataset. The CNN utilizes the same custom framework used to categorize the AFM images based on morphology. The CNN demonstrated similar performance to the previous model, with a testing accuracy of 96%. The model’s predictions were further evaluated using SHAP analysis (Fig. 7).68 Similar to the previous CNN, smaller features were identified at early convolutional blocks, building to larger features in subsequent blocks. In this case, the features extracted from blocks 1 and 2 were strongly correlated to the ‘weakly-ordered’ images, such as short, curved lines and dots (Fig. 7 a, b). However, features in block 3 were primarily correlated to the ‘well-ordered’ images, highlighting parallel groups of lines with minimal curvature (Fig. 7 c). Although the dataset was augmented to balance the ‘weakly-ordered’ class, the model still slightly over-predicted for the ‘well-ordered’ class (Fig. 7 d), similar to the RFC model discussed previously.
Figure 7: SHAP explanations of predictions at convolutional blocks a) 1, b) 2, and c) 3. Positive SHAP values (red) indicate a positive correlation between the highlighted feature and the predicted output, while negative SHAP values (blue) indicate a negative correlation. d) Confusion matrix showing the CNN’s testing performance.
@@ 3.3 Processing-structure relationship prediction
Four traditional ML models, including RF, SVM, MLP, and XGB, were trained on the dataset to predict specific morphological properties of the BCP films as a function of the processing conditions. Training was performed with a 5-fold cross-validation and grid-search, using 80% of the data for training and 20% for testing. The set of hyperparameters evaluated within the grid-search can be found in the SI. The evaluation metric used was the coefficient of determination ($R^{2}$) in which the model with the best average cross-validation $R^{2}$ was chosen to ensure generalizability (Fig. 8).
@@
Figure 8: Parity plots with relative error insets for regression model predictions: a) GISAXS-measured domain spacing and b) FWHM predicted by the RF models, and c) AFM-measured domain spacing and d) grain size predicted by the MLP and SVM models respectively.
@@
RF performed the best for predicting GISAXS-measured domain spacing, achieving a testing $R^{2}$ of 0.80, with the majority of prediction errors falling within ±2.5% (Fig. 8 a). The relatively high accuracy is expected given the low variability and narrow distribution of the domain spacing data measured by GISAXS. However, the model performs much better in the high domain spacing region, around 36nm, where the majority of the data is located, than in the low domain spacing region where data is much more sparse. Similarly, RF outperformed the other models when predicting the FWHM with a testing $R^{2}$ of 0.75 with the majority of error falling within ±25% (Fig. 8 b). The model performs exceptionally well when predicting FWHM within the ‘narrow-FWHM’ regime, around 0.0025 A-1, but struggled with predictions in the ‘broad-FWHM’ region, where data was not as available. Despite the strong correlation and general agreement between the GISAXS and AFM domain spacing measurements, the models trained on domain spacing measured by AFM did not perform as well compared to the models trained on domain spacing measured by GISAXS. XGB performed the best with a testing $R^{2}$ of 0.54, with the majority of error distributed within ±5% (Fig. 8 c). Despite the better distribution of data relative to the GISAXS-measured domain spacing, the high variability of the data likely affected the model’s ability to accurately predict the measurements. Interestingly, SVM was the best performing model for predicting the grain size with a testing $R^{2}$ of 0.32, but with a much more significant error spread than the previous models (Fig. 8 d). The relatively poor accuracy is again likely due to the local nature of the AFM measurements being highly variable across a given sample. Overall, models trained to predict GISAXS-measured properties outperformed those trained on AFM-measured properties, likely because GISAXS captures bulk characteristics, while AFM probes more localized surface features, leading to increased variability within the data. Additional results from the models can be found in the SI.
@@
The RF models trained on GISAXS-measured domain spacing and FWHM were further evaluated with SHAP given their relatively high degree of accuracy to interpret the model’s predictions and evaluate feature importance.70 The mean SHAP value for each processing parameter reflects its overall influence on the model’s predictions. Meanwhile, local SHAP values reflect the influence of a given parameter on the model’s prediction on a per sample basis. In this case, a positive SHAP value indicates that the feature tends to increase the predicted outcome relative to the average prediction. Conversely, a negative SHAP value suggests that the feature generally lowers the prediction relative to the average.
@@
For the RF model trained on GISAXS-measured domain spacing data, additive ratio has a much larger mean SHAP value, 0.95, than solvent ratio, 0.29, and additive type, 0.17 (Fig. 9 a left). This indicates that across all samples, additive ratio has a nearly 3 times greater influence on the resulting BCP domain spacing than the other processing parameters. This is because that the low-volatility additive extends the window for the BCP to rearrange into thermodynamically favorable morphologies, which otherwise remain kinetically trapped. Due to the relatively higher volatility of the solvents, their effects on domain spacing are marginal relative to additive ratio, which is reflected by the substantially lower mean SHAP values. Interestingly, despite small changes in volatility between the additives, the additive type has only a minor overall influence on the predictions, as indicated by its low mean SHAP value.
@@
Furthermore, the local SHAP values were plotted with respect to processing conditions to identify potential cross-relationships between the parameters (Fig. 9 b, c). At low additive ratios, low values of solvent ratio generally lead to slightly higher predicted domain spacing, corresponding to well-ordered samples (Fig. 9 b). This suggests co-solvent systems, solubilizing each segment of the BCP result in better ordered BCP morphologies at low additive ratios. Additionally, as shown in Fig. 9 c, the SHAP values for additive type reveal that the influence of CN on predictions of domain spacing are narrowly aggregated, indicating CN provides good control over BCP morphology. On the other hand, the SHAP values for MN are much more dispersed, and slightly larger than those for CN. Given the slight increase in volatility of MN relative to CN, the increased mobility of the chains does not last as long for MN, therefore kinetically trapping the BCP chains in potentially unfavorable arrangements, leading to more variable, unpredictable domain spacings.

### v2

@@ 2.2.2 AFM
The BCP thin films in this study are expected to arrange in hexagonal cylinder morphologies, given the PS-b-PEO block ratio. However, the orientation and ordering of the cylinders is highly dependent on the processing conditions of the BCP films. hao_self-assembly_2017 Therefore, the 202 AFM images collected represent various orientations of hexagonally packed cylinders, including vertical, horizontal, and a coexistence of both orientations, which manifest as ‘dot’-, ‘line’-, and ‘mixed’-type features in the AFM images, respectively. While no single approach is sufficient to analyze the local ordering, or grain size, of the microdomains for different orientations, multiple approaches can be used in tandem to create a robust AFM analysis approach (Fig. 3). Furthermore, this approach is not limited to characterizing hexagonally packed cylinder morphology, as is performed in this study, and could be extended to other morphologies that similarly yield ’dot’- and ’line’-type surface features.
@@
Figure 3: Overview of the morphology agnostic, feature-based AFM image analysis process for extracting grain size. a) AFM images are classified based on domain orientation by a convolutional neural network (CNN). b) Images containing mixed-type features are segmented into dot- and line-type features. c) Color-wheel analysis of line-type images, color-coding microdomains based on orientation, then segmenting the color-coded microdomains before determining grains. d) Voronoi analysis of dot-type images, features are first isolated before applying the Voronoi analysis and identifying grains.
@@
AFM images are first classified into three categories (Fig. 3 a), in this study depending on the orientation of the hexagonal cylinders, with the use of a convolutional neural network (CNN), discussed in more detail in the following section. After classification, the binarized dot- and line-type images were ready for analysis, however, the mixed-type images required further pre-processing. Segmentation of the mixed-type images was performed using the labelme software to create two distinct masks for the dot- and line-type features (Fig. 3 b). wada_wkentarolabelme_2021 The segmented masks were treated the same as the the dot- and line-type images for further analysis.
[…]
@@
[…]
@@ 2.2.3 Convolutional Neural Network AFM Image Classification
A CNN was first employed to classify the AFM images based on their surface features, specifically, line-, dot-, and mixed-type features, corresponding to vertically aligned cylinders, horizontally aligned hexagonal cylinders, and a combination of both. The CNN was trained using the binarized images as inputs and manually labeled classes as targets. Additionally, due to a significant imbalance of images with line-type features relative to the images with dot- and mixed-type features, the images within the latter classes were augmented, creating three copies of each image rotated in increments of 90°, creating a more balanced dataset. Furthermore, the dot-type images were reflected horizontally and vertically to further increase the amount of data for the class. After augmentation, the dataset was much more balanced, with 171, 84, and 72 images for line-, dot-, and mixed-type features, respectively.
@@
The model follows a simplified VGG-like structure,simonyan_very_2015 consisting of three convolutional blocks with two convolutional layers each, followed by batch normalization, ReLU activations, and max pooling (Fig. 4 a). A more detailed description of the CNN can be found in the SI. Among the 327 AFM images used in the classification task, 228 were allocated for training, 65 for validation, and 34 for testing. On the test set, the CNN model achieved a high classification accuracy of 97.06%. To interpret the model’s decision-making process, the Shapley Additive exPlanations (SHAP) Python package was employed (Fig. 4 b-d). SHAP analysis provides a unified framework for explaining ML model predictions by assigning each feature an importance value, its SHAP value, based on its contribution to the model’s output.lundberg_unified_2017 When SHAP analysis is applied to a CNN trained on images, each pixel (or small region of pixels) is assigned a value that reflects how much it contributed to the prediction, relative to a reference baseline. Regions with positive contributions are shown in red, indicating they increase confidence in the predicted class, while regions with negative contributions are shown in blue, indicating they decrease confidence. In practice, this means that the SHAP maps highlight the specific parts of an image the CNN relies on. For example, in dot-type images, the dot clusters appear as red regions, showing that they are the main features driving the classification.
@@
Figure 4: a) Simplified architecture of the CNN used for surface feature classification. AFM images are passed through a series of convolutional blocks to extract hierarchical feature maps, which are then fed into fully connected layers to produce a probabilistic prediction over the three feature-based classes (line, dot, and mixed). b-d) SHAP-based interpretation of the CNN’s predictions at convolutional blocks 1-3. Positive SHAP values (red) highlight features that positively contribute to the predicted class, while negative SHAP values (blue) indicate features that negatively influence the prediction. e) Confusion matrix summarizing the CNN’s performance on the test set.
@@
Within each convolutional block, features were extracted from the input to generate feature maps. After each block, spatial dimensions were reduced, causing fine-scale features to be captured in the earlier blocks, while progressively larger and more abstract features were analyzed in subsequent blocks. The SHAP analysis of block 1 reveals that features evaluated at smaller length scales exhibit strong positive correlations with dot- and mixed-type classifications, such as dots and disconnections in the line features. This aligns with the physical interpretation of the images, where these features characteristically lack prominent large structural features. Similarly, the SHAP analysis of block 2 reveals positive correlation between clusters of dots with the dot-type classification, and groups and lines with heavy curvature with the mixed-type classification. Interestingly, block 2 also begins to pick up parallel groups of line features, positively correlating them with the line-type classification. Finally, features extracted from block 3 primarily correlate positively with the line-type classification, focusing on long stretches of line features. Overall, the CNN performed exceptionally well distinguishing between dot- and line-type classifications, however, struggled with the mixed-type classifications, over-predicting the dot-type classification for some mixed-type images (Fig. 4 e).
@@ 3.1 Dataset overview
As mentioned previously, the processing parameters explored in this study include solvent ratios of toluene:THF from 0.5 to 1.0, additive ratios from 1.00 to 7.00, and additive types of CN and MN. Domain spacing and FWHM from GISAXS measurements show narrow distributions, with domain spacing measurements around 36 nm and FWHM around 0.0025 A-1 (Fig. 5 a, b). These distributions indicate the majority of the BCP films have relatively good ordering, indicated by the low FWHM value which is reflective of a narrow distribution of domain spacings in the sample. However, the AFM-measured morphological properties, domain spacing and grain size, demonstrate broader distributions (Fig. 5 c, d). While the range of values for the GISAXS- and AFM-based measurements are in agreement with each other, the localized nature of the AFM measurements leads to higher variability, in contrast to GISAXS measurements which reflect the bulk properties of the sample.
@@
Pearson correlation coefficients were calculated to assess potential linear relationships between the processing parameters and the resulting morphological properties (Fig. 5 e). Among all processing parameters, the additive ratio exhibited the strongest Pearson correlation with the measured morphological properties. A previous study using in situ ellipsometry showed that it takes a few minutes for the high boiling point additive to evaporate from the BCP films.ogieglo_situ_2015 During this period, the additive-infused BCP film exhibits enhanced segmental mobility, allowing polymer chains to rearrange into more ordered and thermodynamically favorable morphologies. Interestingly, the type of additive exhibits a relatively strong Pearson correlation with both the GISAXS-measured FWHM and AFM-measured domain spacing. For AFM, this may be attributed to surface-specific effects introduced by the additive, such as differential evaporation dynamics, not captured by bulk GISAXS measurements. In the case of GISAXS-measured FWHM, the correlation may reflect broader distributions of domain spacings within the sample, suggesting increased structural heterogeneity without significantly altering the average domain spacing. Meanwhile, solvent ratio has a relatively low Pearson correlation across all measured outputs. Although solvent plays a critical role in the initial solubilization of the polymers, it has little effect on the resulting morphology of the BCP film, due to the low boiling point solvent volatilizing during the spin coating process. Furthermore, the Pearson correlations were evaluated for the morphological outputs with respect to each other (Fig. 5 f). The domain spacing from both GISAXS and AFM has a fairly large Pearson correlation, indicating good agreement between the measurements from the two approaches. Domain spacing from GISAXS and AFM are inversely correlated with GISAXS-measured FWHM, indicating that as the samples become more ordered, i.e., decreasing FWHM, domain spacing increases. Additionally, the grain size has positive Pearson correlations with domain spacing and an inverse correlation with FWHM, again indicating that as the samples become more ordered, the grain size increases. Scatter plots of the morphological properties are shown in the SI (Fig. S4).
3.2 Domain orientation classification
@@
Classification models were trained using solvent ratio, additive ratio, and additive type as input features, with ‘dot-’, ‘line-’, and ‘mixed-type’ classifications, based on AFM images, as the target output. This approach allows for the creation of a domain orientation map as a function of the processing conditions used to create the sample. Support Vector Classifier (SVC) and Random Forest Classifier (RFC) were selected for this task due to their robust performance and high accuracy in classification problems. The data set was split using 80% as the training set, and 20% for the test set to evaluate the model performance while using a 5-fold cross validation to ensure generalizability. A grid search approach was employed to evaluate various combinations of hyperparameters, which can be found in the SI for each model. The PyTorch accuracy metric was used to evaluate the models, in which the model with the best cross-validation accuracy was selected.
Both classifiers performed exceedingly well, with RFC performing slightly better than SVC, with cross-validation accuracies of 96% and 93%, respectively. Although the RFC model outperformed in overall accuracy, its tree-based architecture yields rigid, stepwise phase boundaries due to poor interpolation between data points. In contrast, SVC defines smoother, continuous boundaries by maximizing inter-class margins, making it more suitable for constructing phase maps. Consequently, SVC was chosen for further analysis. The SVC model overpredicted the line- and mixed-type surface features (Fig. 6 a), likely because the dataset was skewed towards line- (163) and mixed-type (32) features with few samples exhibiting dot-type (7). Two distinct phase maps were generated: one for samples with CN as the additive and another for MN. The CN map showed no misclassifications (Fig. 6 b, left), as the phase transitions were largely dependent on the additive ratio. In contrast, the MN map was more complex, with three misclassifications near phase boundaries (Fig. 6 b, right), likely a consequence of MN’s lower boiling point and greater impact on the solvent ratio. Despite these discrepancies, both systems exhibited a clear trend in which increasing additive ratio promotes a transition from dot-type (vertically stacked cylinders) to line-type (horizontally aligned cylinders) as the kinetic window is extended, allowing the samples to approach their thermodynamic equilibrium state.
Figure 6: a) Confusion matrix showing the testing results of the SVC model. b) Decision boundary plots illustrating the SVC decision boundaries as the transparent regions and the actual domain orientation types as dots. Separate plots were created for the CN additive, left, and the MN additive, right, as a function of solvent and additive ratios. For both additive types, increasing additive ratio promotes a transition from dot-type surface features to line-type surface features, i.e. a transition from vertically oriented domains to horizontally oriented domains as the sample approaches thermodynamic equilibrium.
@@ 3.3 Processing-structure relationship prediction
Four traditional ML models, including RF, SVM, MLP, and XGB, were trained on the dataset to predict specific morphological properties of the BCP films as a function of the processing conditions. The performances of the ML models were compared with more straightforward approaches, including linear, lasso, and Bayesian ridge regression. Results indicated that the ML models consistently outperformed the other approaches across all target properties (Fig. S5), discussed in more detail in the SI. Training was performed with a 5-fold cross-validation and grid-search, using 80% of the data for training and 20% for testing. The set of hyperparameters evaluated within the grid-search can be found in the SI. The evaluation metric used was the coefficient of determination ($R^{2}$) in which the model with the best average cross-validation $R^{2}$ was chosen to ensure generalizability (Fig. 7). Furthermore, the effect of training set size was evaluated to ensure models could not be trained to achieve comparable performance on smaller datasets (Fig. S6).
@@
Figure 7: Parity plots with relative error insets for regression model predictions: a) GISAXS-measured domain spacing and b) FWHM predicted by the RF models, and c) AFM-measured domain spacing and d) grain size predicted by the XGB and SVM models respectively.
@@
RF performed the best for predicting GISAXS-measured domain spacing, achieving a testing $R^{2}$ of 0.80, with the majority of prediction errors falling within ±2.5% (Fig. 7 a). The relatively high accuracy is expected given the low variability and narrow distribution of the domain spacing data measured by GISAXS. However, the model performs much better in the high domain spacing region, around 36 nm, where the majority of the data is located, than in the low domain spacing region where data is much more sparse. Similarly, RF outperformed the other models when predicting the FWHM with a testing $R^{2}$ of 0.75 with the majority of error falling within ±25% (Fig. 7 b). The model performs exceptionally well when predicting FWHM within the ‘narrow-FWHM’ regime, around 0.0025 A-1, but struggled with predictions in the ‘broad-FWHM’ region, where data was not as available. Despite the strong correlation and general agreement between the GISAXS and AFM domain spacing measurements, the models trained on domain spacing measured by AFM did not perform as well compared to the models trained on domain spacing measured by GISAXS. XGB performed the best with a testing $R^{2}$ of 0.54, with the majority of error distributed within ±5% (Fig. 7 c). Despite the better distribution of data relative to the GISAXS-measured domain spacing, the high variability of the data likely affected the model’s ability to accurately predict the measurements. Interestingly, SVM was the best performing model for predicting the grain size with a testing $R^{2}$ of 0.32, but with a much more significant error spread than the previous models (Fig. 7 d). The relatively poor accuracy is again likely due to the local nature of the AFM measurements being highly variable across a given sample. Overall, models trained to predict GISAXS-measured properties outperformed those trained on AFM-measured properties, likely because GISAXS captures bulk characteristics, while AFM probes more localized surface features, leading to increased variability within the data. Additional results from each model can be found in the SI (Figs. S7-S10).
@@
The RF models trained on GISAXS-measured domain spacing and FWHM were further evaluated with SHAP given their relatively high degree of accuracy to interpret the model’s predictions and evaluate feature importance.lundberg_local_2020 The mean SHAP value for each processing parameter reflects its average influence on the model’s predictions, e.g. domain spacing or FWHM. Local SHAP values, by contrast, describe the influence of a parameter on an individual sample’s prediction. A positive SHAP value indicates the parameter drives the predicted property higher relative to the model’s average prediction, while a negative SHAP value drives the predicted property to lower values.
@@
For the RF model trained on GISAXS-measured domain spacing data, additive ratio has a much larger mean SHAP value, 0.95, than solvent ratio, 0.29, and additive type, 0.17 (Fig. 8 a left). This indicates that across all samples, additive ratio has an influence greater than 3 times that of the other processing parameters on the resulting BCP domain spacing. This is because the low-volatility additive extends the temporal window for the BCP to rearrange into thermodynamically favorable morphologies, which otherwise remain kinetically trapped. Due to the relatively high volatility of the solvents, their effects on domain spacing are marginal relative to additive ratio, which is reflected by the substantially lower mean SHAP values. Interestingly, despite small changes in volatility between the additives, the additive type has only a minor overall influence on the predictions, as indicated by its low mean SHAP value.
@@
Furthermore, the local SHAP values were plotted with respect to processing conditions to identify potential cross-relationships between the parameters (Fig. 8 b, c and Fig. S11). At low additive ratios, low values of solvent ratio generally lead to slightly higher predicted domain spacing, corresponding to well-ordered samples (Fig. 8 b). This suggests co-solvent systems, solubilizing each segment of the BCP result in better ordered BCP morphologies at low additive ratios. Additionally, as shown in Fig. 8 c, the SHAP values for additive type reveal that the influence of CN on predictions of domain spacing are narrowly aggregated, indicating CN provides good control over BCP morphology. On the other hand, the SHAP values for MN are much more dispersed, and slightly larger than those for CN. Given the slight increase in volatility of MN relative to CN, the increased mobility of the chains does not last as long for MN, therefore kinetically trapping the BCP chains in potentially unfavorable arrangements, leading to more variable, unpredictable domain spacings.